OpenStudy (anonymous):

what is the work for the distancr between two parallel line: y=2x and y=2x+3 : square root of 18

OpenStudy (anonymous):



OpenStudy (campbell_st):

well you now a point on the line y = 2x + 3.... the y-intercept (0, 3) so you can use the perpendicular distance from a line to a point formula \[d = \frac{|ax_{1} + by_{1} + c|}{\sqrt{a^2 + b^2}}\] so rewrite y = 2x in standard form and use the point (0, 3)